The first section

Second Beamer Example

Till Tantau

Fakultät für Elektrotechnik und Informatik

Technical University of Berlin

April 21, 2006

Till Tantau  TU Berlin

Second Beamer Example The first section

This is a frame with two overlays.

▶ The first item…

▶ … and the second one.

The first section

This is a frame with two overlays.

▶ The first item…

▶ … and the second one.